# Gaussian Splatting Visual MPC for Granular Media Manipulation

### v1

@@ Abstract
Recent advancements in learned 3D representations have enabled significant progress in solving complex robotic manipulation tasks, particularly for rigid-body objects. However, manipulating granular materials such as beans, nuts, and rice, remains challenging due to the intricate physics of particle interactions, high-dimensional and partially observable state, inability to visually track individual particles in a pile, and the computational demands of accurate dynamics prediction. Current deep latent dynamics models often struggle to generalize in granular material manipulation due to a lack of inductive biases. In this work, we propose a novel approach that learns a visual dynamics model over Gaussian splatting representations of scenes and leverages this model for manipulating granular media via Model-Predictive Control. Our method enables efficient optimization for complex manipulation tasks on piles of granular media. We evaluate our approach in both simulated and real-world settings, demonstrating its ability to solve unseen planning tasks and generalize to new environments in a zero-shot transfer. We also show significant prediction and manipulation performance improvements compared to existing granular media manipulation methods.
@@ I Introduction
Neural rendering and view synthesis methods [1, 2, 3, 4, 5, 6] have enabled a wide set of applications in scene understanding, 3D reconstruction and representation. Moreover, they have shown promise in many complex robot manipulation tasks on rigid body objects [7, 8]. Manipulating granular materials such as beans, nuts, rice, oats, and other such objects common in daily life remains a challenging problem, so in this paper we address the question of whether neural rendering methods, Gaussian Splatting in particular, provide a good representation for control of granular media.
Figure 1: Our method takes a few multi-view images of a scene and their corresponding camera poses as input (a) converts them into their Gaussian splatting representation, (b) learns a dynamics model over these representations, and (c) performs visual model-predictive control for granular material manipulation, which requires view synthesis and dynamics rollouts.
Several factors contribute to the difficulty of granular material manipulation. First, modeling the interactions between particles is complicated due to the intricate physics involved [9] and the unknown geometry of individual particles. Second, accounting for all particles in planning requires a high-dimensional state [10, 11], which creates challenges for downstream policy learning or planning algorithms. Third, visually identifying and tracking individual granular particles in a pile is nearly impossible due to their self-similarity, which leads to data association problems, and due to the inherent partial observability of the setting. This has led to the majority of existing works that simulate granular materials doing so without ground truth data from the real world [12, 13, 14], relying on simulated data. Finally, accurately predicting particle dynamics is computationally expensive [9, 15, 16].
@@
In this work, we show that Gaussian splatting [4] over video frames provides effective representations for downstream model-predictive control over granular media. Gaussian splatting is an image rendering and reconstruction technique originating in computer graphics that represents a 3D scene as a collection of Gaussians (splats), each of which is centered around points in the scene and is associated with a single color. An image of the scene from a given viewpoint is rendered by projecting each splat into 2D and blending colors where they overlap. This collection of splats that can also be parameterized by position, and rotation of Gaussian provides a smooth and continuous representation in terms of space of the scene, making it particularly well-suited for modeling video sequences.
@@
We evaluate our approach in a variety of pile manipulation tasks in both simulation and real-world settings, and we show that our method outperforms existing baselines both in terms of dynamics prediction and in terms of task performance. Our model also demonstrates the ability to solve previously unseen complex planning tasks. Furthermore, we showcase the generalizability of our method by transferring a trained model to different environments with varying object shapes in a zero-shot setting.  [^1] For more details, code, videos, and the paper’s appendix please refer to our project website: https://weichengtseng.github.io/gs-granular-mani/.
II Related Works
@@
Another prominent line of research involves the use of implicit neural representations, where 3D shapes and environments are encoded as continuous functions rather than discrete points or voxels. Neural Radiance Fields (NeRF) [1, 6] is a notable example of this approach, where scenes are represented as volumetric radiance fields that can be rendered from arbitrary viewpoints. While originally designed for view synthesis, NeRF and its variants have inspired applications in robotic manipulation, especially in scenarios where precise modeling of object geometry and appearance is critical [24, 25].
Particle Dynamics. Inductive biases, particularly object-centric representations, have been widely adopted in learning-based dynamics models. Particle-based representations serve as strong inductive biases for representing deformable objects. In particular, DPI-Net [12] combines a hierarchical particle dynamics model with MPC-based control for deformable object manipulation. However, particle-based approaches face scalability issues as the number of particles increases [26, 27], thereby making them computationally expensive and challenging to use in practical planning tasks.
@@
To address these limitations, recent research has explored data-driven methods that learn granular dynamics from observations [29]. Techniques like neural networks have been employed to approximate material behavior [28, 16, 17], enabling faster predictions during manipulation tasks. However, these approaches often struggle with generalization across different types of granular materials and varying conditions. Besides, these granular material manipulations only represent material in 2D space, which constrains the potential of 3D manipulation tasks. Our approach leverages advanced 3D reconstruction technique, which alleviates this limitation.
Figure 2: (a) The dynamics model f, conditioned on the input action ${\textbf{{u}}_{t}}$, predicts the temporal evolution of the scene representation zt. During planning time, we calculate the task objective $c(\textbf{{Z}}_{T},\textbf{{Z}}_{target})$ and backpropagate the gradients to optimize the action sequence ${\textbf{{u}}_{t}}$ (b) The dynamics model $f$, conditioned on the input action ${\textbf{{u}}_{t}}$, predicts the temporal evolution of the scene representation $\textbf{{Z}}_{t}$. During planning time, we calculate the task objective $c(\textbf{{Z}}_{T},\textbf{{Z}}_{target})$ and backpropagate the gradients to optimize the action sequence ${\textbf{{u}}_{t}}$.
@@ III Preliminaries
The parameters of the Gaussian can be updated by minimizing the $\mathcal{L}_{1}$ loss and structural similarity index measure (SSIM) $\mathcal{L}_{SSIM}$ between the reconstructed image and ground-truth image.
@@
where $\mu_{x}$ and $\mu_{y}$ are pixel mean of $x$ and $y$. $\sigma_{x}$ and $\sigma_{y}$ are pixel variance of $x$ and $y$. $\sigma_{xy}$ is the covariance of $x$ and $y$. $c_{1}$ and $c_{2}$ are constant to stabilize division.
@@ IV-A Problem Formulation
Given multi-view RGBD observations $\textbf{{O}}_{target}=\{\textbf{{o}}^{v},m^{v}\}_{v=1}^{N}$ of the target pattern of the granular material, where $o^{v}$ represents the RGBD image and $m^{v}$ indicates the corresponding camera pose, we would like to manipulate the granular material to minimize the cost measurement $c$ such that the granular material can match the target pattern. We define the following trajectory optimization problem over a horizon $T$
$\textbf{{u}}_{t}=argmin_{\textbf{{u}}_{t}}c(\textbf{{Z}}_{T},\textbf{{Z}}_{target})\$ (6)
@@
where $h$ is the perception module that performs Gaussian splatting and $f()$ is the dynamic model which predicts the representation’s evolution $\textbf{{Z}}_{t+1}$ from the previous representation and action. The optimization aims to find the action sequence $\{\textbf{{u}}_{t}\}$ to minimize the cost function $c(\textbf{{Z}}_{T},\textbf{{Z}}_{target})$. Also, $\textbf{{u}}_{t}\in\mathbb{R}^{4}$ represents the starting position and pushing direction of the end-effector. To enable learning a dynamics model for granular materials, we collect a dataset $D_{\text{RGBD}}=\{(\textbf{{O}}_{t},\textbf{{u}}_{t},\textbf{{O}}_{t+1})\}$ via a simulator of a manipulator interacting with particles.
@@ IV-C Learning a Visual Dynamics Model
Given the dataset $D_{GS}$, we learn a visual dynamics model over Gaussian splatting representations. Our dynamics model $f()$ is formed as a Graph Neural Network (GNN) [30] with iterative message passing that takes the Gaussian splatting reconstruction as input and predicts the translation and rotation for each Gaussian.
@@
$\mathcal{L}_{gaussian}(n,\hat{n})=||\textbf{{g}}-\hat{\textbf{{g}}}||_{2}+\lambda(1-|\textbf{{r}}_{t+1}\cdot\hat{\textbf{{r}}}|)$ (14)
@@ IV-D Planning
We integrate the solution to Equation 6 into a MPC framework using the optimized action sequence in a closed-loop system. At each MPC step, we follow the procedure outlined in Algorithm 1, which first determines the appropriate resolution for representing the environment. Next, a combination of sampling and gradient descent is used to derive the action sequence through trajectory optimization, employing the shooting method. After executing the first action from the sequence in the real world, new observations are obtained, and Algorithm 1 is applied again. This iterative process allows the system to continuously incorporate environmental feedback and adaptively choose the optimal resolution as the task progresses. Further details on the task objectives and MPC hyperparameters are provided in the supplementary materials.
Data: Current observation $\textbf{{O}}_{t}$, target $\textbf{{O}}_{target}$, planning horizon $T$, the dynamics module $f$, and gradient descent iteration $N$
Result: a sequence of action actions $\textbf{{u}}_{0:T-1}$
Get current representation $\textbf{{Z}}_{t}$ from observation $\textbf{{O}}_{t}$;
and target representation $\textbf{{Z}}_{target}$ from $\textbf{{O}}_{target}$;
Sample $K$ action sequences $\textbf{{u}}^{1:K}_{0:T-1}$ ;
for k = 1,…,K do
       for i = 1,…,N do
             for t = 0,…,T-1 do
                   Predict the next step $\textbf{{Z}}_{t+1}=f(\textbf{{Z}}_{t},\textbf{{u}}_{t})$;
             end for
            Compute the task loss $c^{k}=c(\textbf{{Z}}_{T},\textbf{{Z}}_{target})$;
             update the $\textbf{{u}}^{1:K}_{0:T-1}$ with the task loss;
       end for
end for
$k_{opt}=argmin_{k}c^{k}$ ;
return $\textbf{{u}}^{k_{opt}}_{0:T-1}$ ;
Algorithm 1 Planning algorithm
Figure 3: Real-world experiment setup. (a) The robotic manipulator, equipped with a pusher at the end-effector, moves object piles within the workspace. Four calibrated RGBD cameras mounted around the workspace provide visual observations of the environment. (b) The granular materials used in real-world experiments include coffee beans, peanuts, pistachios, and almonds.
Figure 4: Rollout of our dynamics model. We show the rollout predictions of dynamic model in both simulation (left) and real-world data (right). Both of the rollout results show that the dynamics model prediction is accurate for a few steps.
@@ V-A Implementation Details
Simulation Setup. The simulation environment is based on Pybullet [34], adapted from the Ravens [35] framework. Throughout both data generation and evaluation, unless otherwise specified, we use a set of 50 cubic blocks, each measuring 1 cm in size, along with a planar pusher of 5 cm in length.
Physical Robot Setup. The real-world setup features a Franka Panda manipulator [36] and four Intel RealSense cameras (see Appendix for further details). The camera captures a top-down view of the workspace, with images rectified through homographic warping. Color and depth thresholding are then applied to extract the object density field. We directly transfer a model trained in simulation to real-world experiments. To ensure consistency between simulation and reality, we resize the input images to match the workspace size used in the simulation. The pusher has the same width as its simulated counterpart and is attached to the robot’s gripper. Fig. 3 shows the granular materials tested.
@@
  • Dyn-Res[16] constructs dynamic-resolution particle representations of the environment and learns a unified dynamics model using GNNs. This model allows for continuous adjustment of the abstraction level. During testing, the agent adaptively determines the optimal resolution at each MPC step.
  • NFD[28] employs a fully convolutional neural network that operates on a density field-based representation of object piles and pushers. This approach leverages the spatial locality of inter-object interactions and translation equivariance through convolutional operations.
  • NeRF-dy [37] integrates NeRF with time contrastive learning in an autoencoding framework. It learns viewpoint-invariant, 3D-aware scene representations. By constructing a dynamics model over this learned representation space, NeRF-dy enables visuomotor control for challenging manipulation tasks.
  • DVF [17] is a strong dynamics model that uses field-based state representation without inductive biases. DVF only predicts the start and end position of a straight push.
@@
Metrics. In simulation, we perform 100 trials, while for real-world experiments, we conduct 20 trials. We use two metrics to assess the performance of the framework.
@@ VI Limitations
Manipulation Efficiency. Granular material dynamics are inherently complex due to the intricate interactions between particles. To manage this challenge, we limit the robot arm’s movement speed, reducing the intensity of particle interactions and promoting more stable and controllable manipulation. Particle Size Limitation. Our framework is less suited for manipulating very small particles, such as salt, sugar, or rice. This limitation stems from the difficulty in accurately reconstructing such tiny particles using Gaussian splatting, which struggles to maintain precision at smaller scales. We anticipate however that this would be an issue for any vision-based policy. Shortsighted Planning. While our method effectively optimizes for the next best trajectory based on a global cost function, it focuses on short-term decision-making. In more complex scenarios, planning several steps ahead may be necessary to determine truly optimal actions. Future work will explore longer-horizon predictions and the development of more efficient sampling and optimization strategies to facilitate long-term planning.

### v3

@@ Abstract
Recent advancements in learned 3D representations have enabled significant progress in solving complex robotic manipulation tasks, particularly for rigid-body objects. However, manipulating granular materials such as beans, nuts, and rice remains challenging due to the intricate physics of particle interactions, high-dimensional and partially observable state, inability to visually track individual particles in a pile, and the computational demands of accurate dynamics prediction. Current deep latent dynamics models often struggle to generalize in granular material manipulation due to a lack of inductive biases. In this work, we propose a novel approach that learns a visual dynamics model over Gaussian splatting representations of scenes and leverages this model for manipulating granular media via Model-Predictive Control. Our method enables efficient optimization for complex manipulation tasks on piles of granular media. We evaluate our approach in both simulated and real-world settings, demonstrating its ability to solve unseen planning tasks and generalize to new environments in a zero-shot transfer. We also show significant prediction and manipulation performance improvements compared to existing granular media manipulation methods.
@@ I Introduction
Neural rendering and view synthesis methods [1, 2, 3, 4, 5, 6] have enabled a wide set of applications in scene understanding, 3D reconstruction, and representation. Moreover, they have shown promise in many complex robot manipulation tasks on rigid body objects [7, 8]. Manipulating granular materials such as beans, nuts, rice, oats, and other such objects common in daily life remains a challenging problem, so in this paper, we address the question of whether neural rendering methods, Gaussian Splatting in particular, provide a good representation for control of granular media.
Figure 1: Our method takes a few multi-view images of a scene and their corresponding camera poses as input, and (a) converts them into their Gaussian splatting representation, (b) learns a dynamics model $f$ over these representations, and (c) performs visual model-predictive control for granular material manipulation, which requires view synthesis and dynamics rollouts.
Several factors contribute to the difficulty of granular material manipulation. First, modeling the interactions between particles is complicated due to the intricate physics involved [9] and the unknown geometry of individual particles. Second, accounting for all particles in planning requires a high-dimensional state [10, 11], which creates challenges for downstream policy learning or planning algorithms. Third, visually identifying and tracking individual granular particles in a pile is nearly impossible due to their self-similarity, which leads to data association problems, and due to the inherent partial observability of the setting. This has led to the majority of existing works that simulate granular materials without ground truth data from the real world [12, 13, 14], relying on simulated data. Finally, accurately predicting particle dynamics is computationally expensive [9, 15, 16].
@@
In this work, we show that Gaussian splatting [4] over video frames provides effective representations for downstream model-predictive control over granular media. Gaussian splatting is an image rendering and reconstruction technique originating in computer graphics. It represents a 3D scene as a collection of Gaussians (splats), each centered around points in the scene and associated with a single color. From a given viewpoint, an image of the scene is rendered by projecting each splat into 2D and blending colors where they overlap. This collection of splats that can also be parameterized by the position and rotation of Gaussians provides a smooth and continuous representation of the scene, making it particularly well-suited for modeling video sequences.
@@
We evaluate our approach across a range of granular material manipulation tasks in both simulation and real-world settings, demonstrating its superiority over existing baselines in dynamics prediction and task performance. Our model successfully enables solutions of complex planning tasks. Furthermore, we highlight its generalization capability by transferring a trained model to new environments with varying object shapes in a zero-shot setting.  [^1] For more details, code, videos, and the paper’s appendix, please refer to our project website: https://rvl.cs.toronto.edu/gs-granular-media-mpc/.
II Related Work
@@
Another prominent line of research involves the use of implicit neural representations, where 3D shapes and environments are encoded as continuous functions rather than discrete points or voxels. Neural Radiance Fields (NeRF) [1, 6] is a notable example of this approach, where scenes are represented as volumetric radiance fields that can be rendered from arbitrary viewpoints. While originally designed for view synthesis, NeRF and its variants have inspired applications in robotic manipulation, especially in scenarios where precise modeling of object geometry and appearance is critical [24, 25, 26].
Particle Dynamics. Learning-based dynamics models commonly incorporate inductive biases, particularly object-centric representations. Particle-based representations have been shown to have strong inductive biases when representing deformable objects. For instance, DPI-Net [12] combines a hierarchical particle dynamics module with MPC-based control for deformable object manipulation. However, particle-based approaches suffer from scalability issues as the number of particles increases [27, 28], which makes them computationally expensive and impractical in planning tasks.
@@
To address these limitations, recent research has explored data-driven methods that learn granular dynamics from observations [30]. Techniques like neural networks have been employed to approximate material behavior [29, 16, 17], enabling faster predictions during manipulation tasks. However, these approaches often struggle with generalization across different types of granular materials and varying conditions. Besides, these granular material manipulations only represent material in 2D space, which constrains the potential of 3D manipulation tasks. Our approach leverages an advanced 3D reconstruction technique, which alleviates this limitation.
Figure 2: Our framework. (a) Given demonstration trajectories with multi-view observations, we leverage Gaussian splatting representations to reconstruct the observed images at each timestep. (b) The dynamics model $f$ predicts the temporal evolution of the Gaussian Splatting representation $Z_{t}$ with input action ${\textbf{{u}}_{t}}$. In the planning stage, we optimize the action sequence ${\textbf{{u}}_{t}}$ by minimizing the task objective $c(\textbf{{Z}}_{T},\textbf{{Z}}_{target})$.
@@ III Preliminaries
The parameters of the Gaussians can be updated by minimizing the $\mathcal{L}_{1}$ loss and structural similarity index measure (SSIM) $\mathcal{L}_{SSIM}$ between the reconstructed image and ground-truth image.
@@
where $\mu_{x}$ and $\mu_{y}$ represent pixel mean of $x$ and $y$. $\sigma_{x}$ and $\sigma_{y}$ indicate pixel variance of $x$ and $y$. $\sigma_{xy}$ is the covariance of $x$ and $y$, and $c_{1}$ and $c_{2}$ are constant to stabilize division.
@@ IV-A Problem Formulation
Given multi-view RGBD observations $\textbf{{O}}_{target}=\{\textbf{{o}}^{v},m^{v}\}_{v=1}^{N}$ of the target pattern of the granular material, where $o^{v}$ represents the RGBD image and $m^{v}$ indicates the corresponding camera pose, we would like to manipulate the granular material to minimize the cost measurement $c$ such that the granular material can match the target pattern. The trajectory optimization problem over a horizon $T$ can be defined as follows:
$\textbf{{u}}_{0:T-1}=argmin_{\textbf{{u}}_{0:T-1}}c(\textbf{{Z}}_{T},\textbf{{Z}}_{target})\$ (6)
@@
where $h$ is the perception module that performs Gaussian splatting and $f(\cdot)$ is the dynamic model which predicts the representation’s evolution $\textbf{{Z}}_{t+1}$ from the previous representation and action. This optimization process aims to acquire the sequence of actions $\{\textbf{{u}}_{t}\}$ to minimize the cost function $c(\textbf{{Z}}_{T},\textbf{{Z}}_{target})$. Also, $\textbf{{u}}_{t}\in\mathbb{R}^{4}$ represents the starting position and pushing direction of the end-effector. To enable learning a dynamics model for granular materials, we collect a dataset $D_{\text{RGBD}}=\{(\textbf{{O}}_{t},\textbf{{u}}_{t},\textbf{{O}}_{t+1})\}$ via a simulator of a manipulator interacting with particles.
@@ IV-C Learning a Visual Dynamics Model
Given the dataset $D_{GS}$, we learn a visual dynamics model over Gaussian splatting representations. Our dynamics model $f(\cdot)$ is structured as a Graph Neural Network (GNN) [31] with iterative message passing that takes the Gaussian splatting reconstruction as input and predicts the translation and rotation for each Gaussian.
@@
$\mathcal{L}_{gaussian}(\textbf{{n}},\hat{\textbf{{n}}})=||\textbf{{g}}-\hat{\textbf{{g}}}||_{2}+\lambda(1-|\textbf{{r}}_{t+1}\cdot\hat{\textbf{{r}}}|)$ (14)
@@ IV-D Planning
We perform the optimization shown in Equation 6 as part of a gradient-based MPC loop (as shown in Alg. 1) and we execute the optimized action sequence in a closed-loop fashion. After the first action in the sequence is executed by the robot, we re-run the planning algorithm to generate a new action sequence.
Data: Current observation $\textbf{{O}}_{t}$, target $\textbf{{O}}_{target}$, planning horizon $T$, the dynamics model $f$, Number of sampled action sequence $K$ and gradient descent step $N$
Result: a sequence of actions $\textbf{{u}}_{0:T-1}$
1
2Get current representation $\textbf{{Z}}_{t}$ from observation $\textbf{{O}}_{t}$;
3 Acquire target representation $\textbf{{Z}}_{target}$ from $\textbf{{O}}_{target}$;
4
5Sample $K$ initial action sequences $\textbf{{u}}^{1:K}_{0:T-1}$ ;
6
7for k $\leftarrow$ 1 to K do
8       for i $\leftarrow$ 1 to N do
9             for t $\leftarrow$ 0 to T-1 do
10                   Predict next representation $\textbf{{Z}}_{t+1}=f(\textbf{{Z}}_{t},\textbf{{u}}_{t})$;
11
12            Compute the task loss $c^{k}=c(\textbf{{Z}}_{T},\textbf{{Z}}_{target})$ for the action sequence $\textbf{{u}}^{k}_{0:T-1}$ with $k=1...K$;
13             Optimize $\textbf{{u}}^{k}_{0:T-1}$ with $\nabla_{\textbf{{u}}^{k}_{0:T-1}}c^{k}$
14
[…]
16 return $\textbf{{u}}^{k_{opt}}_{0:T-1}$ ;
Algorithm 1 Our Visual MPC Planning Algorithm
Figure 3: Real-world experiment setup. (a) The robotic manipulator, with a pusher attached to the end-effector, moves granular materials within the workspace. Four calibrated RGBD cameras are mounted around the workspace to provide multi-view observations. (b) The granular materials used in real-world experiments include coffee beans, peanuts, pistachios, and almonds.
Figure 4: Dynamics model rollouts. We show the rollout predictions of the dynamics model in both simulation (left) and real-world data (right). Both of the rollout results show that the dynamics model prediction is accurate for a few steps.
@@ V-A Implementation Details
Simulation Setup. The simulation environment is based on PyBullet [35], adapted from the Ravens [36] framework. Throughout both data generation and evaluation, unless otherwise specified, the granular material is formed by a set of 50 cubic blocks, each measuring 1 cm in size, along with a planar pusher of 10 cm in length.
Physical Robot Setup. Our real-world experimental setup includes four Intel RealSense cameras and a Franka Panda manipulator [37]. More details are provided in the Appendix, found on our project website. The camera captures top-down views of the workspace from different angles. We directly transfer our model trained in the simulation environment to our real-world experiment setup. To alleviate the sim-to-real gap, we resize the input images to match the workspace size used in the simulation. Also, we align the width of the pusher, which is attached to the robot’s gripper with its simulated counterpart Fig. 3 shows the granular materials tested and the robot setup.
@@
  • Dyn-Res[16] constructs dynamic-resolution particle representations for granular materials and learns a unified dynamics model with GNN. This approach enables continuous adjustment of the abstraction level, allowing the agent to adaptively select the optimal resolution at each MPC step during testing.
  • NFD[29] uses a fully convolutional neural network built on a density field-based representation of object piles and pushers. This approach leverages the spatial locality of inter-object interactions and translation equivariance through convolutional operations.
  • NeRF-dy [38] leverages NeRF to learn viewpoint-invariant and 3D-aware scene representations. By constructing a dynamics model over this learned representation space, NeRF-dy enables visuomotor control for challenging manipulation tasks, such as pouring liquid.
  • DVF [17] is a dynamics model that uses field-based state representation without inductive biases.
@@
Metrics. In simulation, we perform 100 trials, while for real-world experiments, we conduct 20 trials. We use two metrics to evaluate the frameworks.
@@ VI Limitations
Manipulation Efficiency. Granular material dynamics are inherently complex due to the intricate interactions between particles. To manage this challenge, we limit the robot arm’s movement speed, reducing the intensity of particle interactions and promoting more stable and controllable manipulation.
Particle Size Limitation. Our framework is less suited for manipulating very small particles, such as salt, sugar, or rice. This limitation stems from the difficulty in accurately reconstructing such tiny particles using Gaussian splatting, which struggles to maintain precision at smaller scales. We anticipate however that this would be an issue for any vision-based policy.
Shortsighted Planning. Despite the fact that the proposed approach effectively optimizes for the next best trajectory based on a cost function, it focuses on short-term decision-making. In more complicated tasks, such as more particles or a complex target pile shape, planning several steps ahead may be necessary to determine truly optimal actions.